a stone thrown down with speed u takes a time t1 to reach the ground while another stone thrown upwards from the same point woth the same speed takes time t2 the maximim height the second stone reaches from the ground is
 

Grade:11

1 Answers

Shailendra Kumar Sharma
6 years ago
 lets say the stones were thrown from a height H 

Now, the case of t1 the stone was thrown downward. 
H=ut1+(1/2)gt12---------------(a)
 
for second case -H =ut2+(1/2)gt22 …............(b)
Maximum height with respect to the level H can be calculated by v2=u2-2as => s= u2/2g so with respect to ground the maximum height will be H+u2/2g But H is not given So we have to find that out which can be found by subtraction a and b equations

2H= {ut1+(1/2)gt12} – {ut2+(1/2)gt22} =u(t1-t2)-(1/2)g(t12-t22
H= {u(t1-t2)-(1/2)g(t12-t22)}/2
so the maxium height with respect to groun is ={u(t1-t2)-(1/2)g(t12-t22)}/2+u2/2g